# references.bib

### v1

@@ Abstract
The greedy algorithm based route planning problem is a method of finding the optimal or near optimal route between a given starting and ending point. This article first uses PCA method to reduce the dimensionality of urban evaluation indicators, extracts key principal components, and KMO and TOPSIS algorithms to reduce the dimensionality of the data. Secondly, for datasets that have not passed the KMO test, a comprehensive evaluation will be conducted using the entropy weight method and TOPSIS method. Finally, based on the greedy algorithm, a route planning algorithm was proposed and optimized to provide personalized route customization according to the different needs of tourists. We also took into account the local travel efficiency, the time required to visit tourist attractions, and necessary daily rest time to reduce costs and avoid falling into the local optimal solution.
@@ III Methods
In order to simplify the analysis process and focus on the core elements, we used principal component analysis (PCA) technology to effectively reduce the dimensionality of urban evaluation indicators and extract several key principal components. For some datasets that did not pass the KMO test, we innovatively combined the entropy weight method and TOPSIS method to achieve comprehensive evaluation and weight allocation of the data. When constructing the route optimization model, we applied a greedy algorithm strategy to ensure the optimization and efficiency of the overall tourism planning scheme.
@@
Figure 1: The Results of the TOPSIS and KMO Downscaling
We comprehensively applied the entropy weighting method and TOPSIS evaluation method, two advanced evaluation tools, to conduct an in-depth analysis of multiple dimensions, including city size, ecological environment, cultural and historical depth, transportation accessibility, climate characteristics, and specialties and cuisines. After a rigorous evaluation process, we have carefully selected 50 cities that are highly attractive to international travelers. This selection not only reflects the comprehensive strength of these cities, but also demonstrates their unique charms and characteristics. TOPSIS is a common and effective multi-attribute decision-making method, which is used to rank the advantages and disadvantages of options by calculating the relative distances of each option from the theoretical optimal point and the theoretical disadvantage point. The TOPSIS algorithm can be found in Pseudocode 3.
The final results are shown in Figure 2 below. The figure shows that the highest rated city is about 0.75, and the subsequent cities have closer ratings, indicating that these cities are relatively close in terms of comprehensive indicators.
Figure 2: TOPSIS Score
@@
Finally, we can get the travel planning route based on the greedy algorithm model, see Figure 3. In making the travel planning, we take the tourists’ interest, tour time, and budget into consideration in order to provide the most cost-effective travel plan.

### v2

@@ Abstract
The route planning problem based on the greedy algorithm represents a method of identifying the optimal or near-optimal route between a given start point and end point. In this paper, the PCA method is employed initially to downscale the city evaluation indexes, extract the key principal components, and then downscale the data using the KMO and TOPSIS algorithms, all of which are based on the MindSpore framework. Secondly, for the dataset that does not pass the KMO test, the entropy weight method and TOPSIS method will be employed for comprehensive evaluation. Finally, a route planning algorithm is proposed and optimised based on the greedy algorithm, which provides personalised route customisation according to the different needs of tourists. In addition, the local travelling efficiency, the time required to visit tourist attractions and the necessary daily breaks are considered in order to reduce the cost and avoid falling into the locally optimal solution.
@@ III Methods
In order to streamline the analytical process and concentrate on the fundamental aspects, we employed the principal component analysis (PCA) technique utilising the MindSpore framework. This approach effectively reduced the dimensionality of the urban evaluation indicators and extracted multiple pivotal principal components. In the case of some datasets that did not pass the KMO test, an innovative approach was taken whereby the entropy weight method was combined with the TOPSIS method in order to achieve a comprehensive evaluation and weight allocation of the data. When constructing the route optimisation model, the greedy algorithm strategy was adopted in order to ensure the optimisation and efficiency of the overall tourism planning scheme.
@@
Figure 1: Dimensionality Reduction Results: MindSpore TOPSIS & KMO
The MindSpore framework was employed to conduct an in-depth analysis of a number of factors, including city size, ecological environment, depth of culture and history, accessibility, climatic characteristics, and specialties and cuisines. This was achieved through the utilisation of two advanced evaluation tools, namely the entropy weighting method and the TOPSIS evaluation method. Following a comprehensive evaluation process, 50 cities were identified as being particularly appealing to international travellers. The selection not only reflects the comprehensive strength of these cities, but also demonstrates their unique charms and characteristics. TOPSIS is a commonly used and effective multi-attribute decision-making method, which calculates the relative distances of each option from the theoretical optimal point and the theoretical disadvantage point to rank the options in order of merit. The TOPSIS algorithm is shown in Pseudocode 3.
The final results of the MindSpore framework experiment are presented in the following Figure 2. As can be observed from the figure, the city with the highest score is approximately 0.75, and the scores of the subsequent cities are more closely aligned, indicating that the overall indicators of these cities are relatively similar.
Figure 2: MindSpore TOPSIS Score
@@
At last, the travel planning routes can be generated using the greedy algorithm model within the MindSpore framework (see Figure 3). When devising a travel plan, factors such as tourists’ interests, tour duration and budget will be considered in order to provide the most cost-effective travel plan.